In an isolated chamber 1kg of steam at 200degree celcius is mixed with 1kg of ice at -200 degree Celsius find the equilibrium temperature and mixture content
 

Grade:11

1 Answers

Prity
6 years ago
Ice(-20° C) to ice(0°C) to water(0°C) to water (100°C) to steam(100°C) to steam(200°C). Q1.  Q2.  Q3.  Q4.  Q5.
Q1=ms?t=10k cal ,similarly Q2=80kcal,Q3=100 kcal,Q4=540 kcal,Q5=50kcal 
Let m at Q3 means mass be m at 100°C of water
Then,
50+m*540=190
m=14/54=7/27
m of steam=1-7/27=20/27
m of water=1+7/27=34/37
And temperature is 100°C